A publishing company plans to publish a book. From the sales data of other publisher of similar books, it works out the demand function for the book as:                                               

Q = 5000 – 5P

Find out

i)                  Demand curve

ii)                Number of book sold at P = Rs. 25

iii)              Price for selling 2500 copies

iv)              Price for zero sales

Elasticity for fall in price from Rs. 25 to Rs. 20


1
Expert's answer
2021-06-02T12:33:59-0400

Solution


I.) Demand curve

P=25. Qd=5000-5(25)=4875

P=35. Qd=5000-5(35)=4875

P=50. Qd=5000-5(50)=4750

P=60.Qd=5000-5(60)=4700

P=80. Qd=5000-5(80)=4600



Just as shown from the above calculations the demand curve of the company slopes downwards..this shows an increase in price of the books leads to a decrease in the demand of books.

ii.)

Books sold at P= Rs.25

P=25

Qd=5000-5(25)=4875

The books sold are 4875@ Rs.25

III.)

Price for selling 2500 copies

Q=2500

2500=5000-5P

5P=5000-2500

P=500

iv.)

Price for zero sales

Q=5000-5p

0=5000-5P

5P=5000

P=Rs.1000


Elasticity for fall from Rs 25 to Rs 20


Elasticity of demand using midpoint formula;


"\frac{Q_2-Q_1}{(Q_2+Q_1)/2} \times100"


"\frac{4900-4875}{(4900+4875)/2}\times100=0.5%"


"\frac{P_2-P_1}{(P_2+P_1)/2}\times100"


"\frac{5}{22.5}\times100=22.2"


Elasticity="\frac{0.5}{22.2}=0.0225"